OpenStudy (anonymous):

Currently you have $80 and your sister has $145. You decide to save $6 of your allowance each week, while your sister decides to spend all her allowance plus $7 each week. How long will it be before you have as much money as your sister?

OpenStudy (anonymous):

Y(ou)=80+6*X Sister(S)=146-7x your money will be equal when Y=S or 80+6*x=145-7x that is 13x=65 x=5

OpenStudy (anonymous):

I see, thank you.(:

OpenStudy (anonymous):

np